# SUGAR: A Sweeter Spot for Generative Unlearning of Many Identities

### v1

@@ 1 Introduction
Despite the undeniable benefits of generative AI across various domains such as image synthesis and content creation [liao2022text, zhang2022styleswin, zhao2022generative, lee2022exp], growing concerns have emerged regarding their potential misuse and unintended consequences [zlateva2024conceptual, ferrara2024genai, zucca2025regulating]. In particular, generative models can inadvertently memorize and reproduce identifiable faces or proprietary images from their training data, posing significant risks to intellectual property rights and personal privacy [shokri2017membership, hu2022membership, dealcala2024comprehensive].
@@
These issues have urged the right to be forgotten [rosen2011right, zhang2024right], a movement supporting individuals’ ability to request the removal of personal data (e.g., images of one’s face) from digital systems like AI models. While possible to remove requested images from training sets, retraining models from scratch is computationally impractical, especially if there are multiple requests over time [li2025machine, yao2025large].
To address the growing need for post-hoc data removal, recent work has explored generative unlearning—the task of removing specific information from pretrained generative models without retraining from scratch [bourtoule2021machine, zhang2023review, heng2024selective, seo2024generative]. Approaches vary in how they suppress unwanted identities. Selective Amnesia [heng2024selective] targets diffusion and VAE-based models, using continual learning techniques to push erased concepts into noise-like outputs. GUIDE [seo2024generative], the first to address Generative Identity Unlearning (GUI) in GANs, instead manipulates the latent space by redirecting target embeddings toward a fixed mean identity representation. As illustrated in Figure 1, both methods demonstrate initial effectiveness for single-identity removal but face key limitations: (a) they lack support for forgetting multiple identities at once or over time; (b) their suppression mechanisms—mapping to noise or an average face—may leak information about the forgotten identity; and (c) they often degrade unrelated generations, causing semantic drift in retained identities. For example, in GUIDE, retained faces in the Baseline row of Figure 1 begin to converge toward the median face, reducing output diversity and model utility.
@@ 2 Related Works
Mitigating Misuse of Generative Models. As generative models continue to advance in their ability to synthesize and manipulate highly realistic images, concerns about ethical misuse have intensified. The integration of powerful image editing techniques—such as latent space manipulation [shen2020interfacegan], language-guided editing [patashnik2021styleclip], and photorealistic face retouching [yoo2019photorealistic]—has made it increasingly accessible for users to modify images of real individuals or apply stylistic transformations [somepalli2023diffusion]. More alarmingly, recent studies have shown that these models can be exploited to extract memorized, copyrighted training data [carlini2023extracting], often without the consent of the individuals or content owners involved. One of the most pressing risks is the unauthorized synthesis of a person’s likeness in misleading or harmful contexts, such as deepfakes or defamatory media [abbas2024unmasking, liu2024metacloak, romero2024generative]. To counter this, Thanh et al. [van2023anti] developed an adversarial attack against UNet, effectively corrupting generated images. Other works in the concept erasure domain focus on eliminating certain visual concepts. For example, Gandikota et al. [gandikota2023erasing] were among the first to introduce removing specific concepts from diffusion models using negative guidance. This is still an ongoing research direction with follow-up works such as [gandikota2023erasing, lu2024mace, petsiuk2024concept] to better improve the balancing between unlearning efficiency and maintaining model utility on remaining concepts. Previous methods are identity-specific, requiring tailoring for each individual, which makes them non-scalable for protecting multiple identities and introduces substantial computational overhead. Moreover, these approaches primarily focus on associations between concepts, such as the “Van Gogh style,” and generated images, a focus more relevant to multi-modal learning, like text-to-image models. In contrast, our study centers on image-to-image models.
@@
Generative Unlearning. Machine unlearning is a technique that enables a machine learning model to “forget” or remove the influence of specific data points from its training data, without needing to be retrained from scratch [tiwary2023adapt, bourtoule2021machine, zhang2023review]. The usage of approximate unlearning is more preferable in the context of GenAI because these models are often trained on large corpora and datasets, making the retraining approach infeasible [xu2024machine, li2025single]. The interest in machine unlearning with generative models has grown recently, where the field was dominated by supervised unlearning before, i.e., classification tasks [gupta2021adaptive, tarun2023fast, golatkar2020eternal, golatkar2021mixed, chundawat2023zero, golatkar2020forgetting]. Seo et al. [seo2024generative] were the first to address GIU for GAN-based models in an approach called GUIDE. However, their method does not account for practical scenarios involving (i) simultaneous unlearning requests—an essential requirement in real-world applications where multiple individuals may request data removal at once—or (ii) sequential unlearning, where requests arrive at different times [zhang2024unlearncanvas]. Our empirical results also demonstrate that while GUIDE is effective for unlearning, it hampers the model’s ability to maintain overall generation quality on retained identities.
@@
Recent studies have also addressed generative unlearning by mapping forgotten identities to a Gaussian distribution $\mathcal{N}(0,\sigma)$ while preserving the original distribution for remaining data in an approach called Selective Amnesia (SA) [heng2024selective, feng2025controllable]. However, using Gaussian noise to replace forgotten identities is suboptimal, as it fails to accurately represent certain training data distributions, potentially corrupting unforgotten identities due to noise interference. Furthermore, SA, as well as GUIDE, (iii) raise security concerns, as the association between a forgotten identity and its output (noise or average face) can still be detected. To this end, we introduce SUGAR to address each of these concerns (i, ii, and iii).
@@ 3.1 Problem Formulation
Problem formulation (multi-identity unlearning). We build on the GUI setup from Seo et al. [seo2024generative] and extend it to handle multiple identity unlearning. We consider a pretrained EG3D-style pipeline $(E_{\psi},G_{\theta},R)$ and aim to obtain an unlearned generator $G_{\theta}^{u}$ that forgets a set of $K$ target identities while retaining performance on others. Let $\mathcal{D}=\mathcal{D}_{u}\cup\mathcal{D}_{r}$ denote the full dataset, where $\mathcal{D}_{u}$ contains images of the $K$ identities to remove and $\mathcal{D}_{r}$ the remainder ($|\mathcal{D}_{r}|\!\gg\!|\mathcal{D}_{u}|$). We operate in latent space: $W_{u}:=\{\,w=E_{\psi}(x)\mid x\in\mathcal{D}_{u}\,\}$ and $W_{r}:=\{\,w=E_{\psi}(x)\mid x\in\mathcal{D}_{r}\,\}$. We keep the mapping network and renderer $R$ fixed and update only $G$.
@@
Remapping function. To disassociate image generation from an identity $\mathcal{I}_{u}$ with latent vector $w_{u}$, a straightforward and effective strategy—used in GUIDE [seo2024generative]—is to remap $w_{u}$ to a surrogate target identity $I^{avg}$ (e.g., the median face) with latent vector $w_{avg}$. With GUI, we adopt this idea and denote the GUIDE loss as the remapping objective.
@@ 3.2 ID-specific De-indentification
In contrast to existing approaches in the unlearning literature—where the target “concepts” for forgotten entities are predefined or manually selected—we propose a learnable and sample-specific procedure to determine where a forgotten sample should be mapped to. We refer to this mechanism as the ID-specific De-identification process.
Following Seo et al. [seo2024generative], we define the identity latent vector $w_{id}$ for an individual as the difference between their latent representation $w_{u}$ and the average latent vector $\bar{w}$, i.e., $w_{id}=w_{u}-\bar{w}$. In our method, we train a mapping model $\mathcal{T}_{\Theta}$ to produce a new identity vector $w_{id}^{t}$ given any input latent vector $w_{u}\in\mathbb{R}^{d\times h\times w}$, such that:
@@
The intuition is that we learn a mapping $\mathcal{T}_{\Theta}$ that assigns each forgotten identity $w_{u}\in W_{u}$ a personalized counterfactual $w_{t}$: a different identity that (i) preserves coarse facial attributes of $w_{u}$ while (ii) being recognizably distinct by facial recognition metrics such as ID or LPIPS [seo2024generative]. In latent space, $w_{t}$ serves as the surrogate for $w_{u}$; after unlearning, $G_{\theta}^{u}$ synthesizes faces consistent with $w_{t}$ rather than $w_{u}$. To learn this identity replacement, we train a transformation function $\Theta$ that maps each $w_{u}\in W_{u}$ to a replacement latent code $w_{t}=\Theta(\cdot,w_{u})$ that meets the above criteria. The transformation is supervised using the following objective:
@@ 3.3 Forgetting in a Continual Learning Fashion
Enforcing Identity Forgetting. Building on the remapping objective of Seo et al. [seo2024generative] in Equation 1, we replace the fixed median-face target with a learnable de-identification target produced by our operator $\Theta(\cdot)$; that is, the average face $\mathcal{I}^{avg}$ is replaced by $\Theta(w)$. As noted in the GUIDE baseline, successful unlearning must also suppress identity information in the local neighborhood of each forgotten latent. We therefore sample a perturbed neighbor $w_{u,a}$ for each forgotten code $w_{u}$:
@@
Elastic Weight Consolidation (EWC). To efficiently preserve behavior in regions most susceptible to collateral damage, we adopt Elastic Weight Consolidation (EWC) [kirkpatrick2017overcoming]. We treat the vicinity probe set $\widetilde{W}$ as a proxy for prior behavior to be retained and regularize parameters that are important for agreement with the source model around forgotten latents.
@@ 4.1 Experimental Setup
Datasets. Following the settings of Seo et al. [seo2024generative], we evaluated GUIDE in three forgetting scenarios: Random, InD (in-domain), and OOD (out-of-domain). These scenarios represent different sources of identities to unlearn. Specifically, identities are sampled randomly from the latent space (Random), from the FFHQ dataset [karras2019style] used for pre-training (InD), or from the CelebA-HQ dataset [karras2018progressive] as unlearning targets (OOD). For both in-distribution (InD) and out-of-distribution (OOD) settings, we use a GAN inversion network $\mathcal{I}$ to map each forgetting image $x_{u}^{i}\in\mathcal{D}_{u}$ to a latent code $w_{u}^{i}=E_{\psi}(x_{u}^{i})$, forming the forget set $W_{u}=\{w_{u}^{i}\}_{i=1}^{N}$; the retaining set $W_{r}$ is sampled from the same datasets and inverted analogously.
Baselines. We select GUIDE [seo2024generative] as our baseline for comparison, as it aligns with our focus on identity-specific unlearning. GUIDE effectively removes identity-related features while preserving the overall structure and quality of the generated images. In contrast, other methods, such as Selective Amnesia (SA) [heng2024selective] and Feng et al. [feng2025controllable], map forgotten identities to Gaussian noise. This approach often results in unnatural artifacts and disrupts facial reconstruction, making it unsuitable for our task. Specifically, mapping to Gaussian noise prevents the generation of coherent and realistic reconstructions after unlearning, which is a critical requirement in our setting.
Evaluation Metrics. We use Identity Similarity (ID), computed with the face recognition network CurricularFace [huang2020curricularface], to measure the similarity between images of faces generated from the same latent codes before and after unlearning. Additionally, we use the Fréchet Inception Distance (FID) score [heusel2017gans] to quantify the overall visual quality and distribution alignment of the generated images. These metrics are calculated on both the forgetting set $\mathcal{D}_{u}$ and the retaining set $\mathcal{D}_{r}$, where $\mathcal{D}_{r}$ is sampled from the same datasets to which $\mathcal{D}_{u}$ belongs.
@@ 4.2.1 Balancing Forgetting and Retaining
#IDS Methods
FFHQ
(In-domain Distribution)
CelebAHQ
(OOD Distribution)
Random
@@
ID GUIDE
0.2773
      $\pm$0.0099
0.0040
      $\pm$0.0077
0.0095
      $\pm$0.0017
-0.0275
      $\pm$0.0034
-0.0431
      $\pm$0.0018
Ours
0.3576
      $\pm$0.0146
0.2664
      $\pm$0.0134
0.2559
      $\pm$0.0078
0.2917
      $\pm$0.0050
0.2433
      $\pm$0.0051
FID GUIDE
133.90
      $\pm$6.9372
177.27
      $\pm$2.2242
163.83
      $\pm$0.9667
143.44
      $\pm$0.8456
146.43
      $\pm$0.6706
Ours
115.98
      $\pm$7.7846
125.98
      $\pm$0.8796
125.93
      $\pm$0.3963
105.40
      $\pm$0.6074
106.12
      $\pm$1.4283
@@ 5 Ablation Study
$d$-Controllable Unlearning.
Our method can be used as a controllable unlearning solution in I2I generative models, where $d$ is the control coefficient. In Equation 3, with a well-trained model $\Phi$, the targeted identity for each identity $w_{id}$ is adjusted with a different coefficient $d$. As shown in Figure 6, the higher the value of $d$, the more features of the forgetting ID $w_{id}$ are retained in the new mapping identity $w_{t}$. With our method, the model vendor can select this controllable parameter to balance the trade-off and the forgetting level that reaches a consensus between the vendor and the requested unlearning party. However, there is a trade-off between forgetting and retaining utility, as discussed in previous work [feng2025controllable, zhang2023review]. This relationship is presented in Table 3. While reducing the distance can make the model forget the forgetting set $\mathcal{D}_{f}$ more drastically (i.e., lower ID and higher FID for the forgetting set), this leads to a greater impact on the retaining set (i.e., lower ID). It is worth noting that even with $d=30$, the generated images are much different from the original identities.
@@
Dual Objective: Identity Removal and Global Utility Preservation.
We study the impact of our two loss terms: forgetting neighbor loss $\mathcal{L}_{nei}$ for thorough identity forgetting and the global preservation loss $\mathcal{L}_{\text{ewc}}$ for utility retention. To evaluate identity removal, we conduct a multi-image test on CelebAHQ identities, assessing whether our method forgets other images of the same identity that are not explicitly included in $W_{u}$. Figure 9 summarizes the results. “Ours-v1” denotes the variant without the neighbor loss $\mathcal{L}_{nei}$ in Equation 5, so the forgetting objective applies only to the anchor images of each target identity. Without this term, residual attributes of the forgotten identity can persist, even when the generated images depict distinct individuals. Adding $\mathcal{L}_{nei}$ enforces consistency across the local neighborhood in latent space, which is crucial for the ultimate goal of removing an entire identity rather than only a single image. To analyze $\mathcal{L}_{ewc}$, Figure 9 compares reconstructions with (Ours) and without (Ours-v2) the global preservation loss. The first two retained identities come from the same distribution as the forgotten ones, while the last three (IDs 3–5) are randomly sampled latents. Random identities remain largely unaffected by unlearning (even under GUIDE), whereas closer identities (IDs 1–2) are more susceptible to collateral forgetting. Ours-v2 already preserves these better than GUIDE due to our de-identification, and adding $\mathcal{L}_{ewc}$ further improves retention by preserving fine details (e.g., glasses, hair, wrinkles). We provide a detailed quantitative result for our ablation study in Appendices B.1. and B.2.
Sequential unlearning.
We consider the setting where unlearning requests arrive sequentially rather than in a single batch. Let $G_{\theta}^{(0)}\!=\!G_{\theta}^{*}$ be the pretrained model and let $\mathcal{D}_{f}^{(k)}$ denote the (new) identities to forget at stage $k{=}1,\dots,4$. At each stage, we initialize from the previously unlearned model and run the same unlearning procedure (architecture, losses, schedule, and hyperparameters unchanged): $G_{\theta}^{(k)}\;\leftarrow\;\textsc{Unlearn}\!\big(G_{\theta}^{(k-1)},\,\mathcal{D}_{f}^{(k)}\big).$ Figure 9 shows qualitative results after four stages for GUIDE and SUGAR. While GUIDE accumulates collateral damage—progressively eroding features of the retained identities across stages—SUGAR continues to forget the requested identities while preserving the appearance of the retained set. Detailed setup and additional analysis are provided in Appendix B.5.
@@
Forgetting
GUIDE
254.31
$\pm$ 376.36
25.4262
$\pm$ 0.2026
0.2048
$\pm$ 0.0000
25.4262
$\pm$ 0.2026
Ours
1022.81
55301.1
18.5459
$\pm$ 0.8418
0.2936
$\pm$ 0.0002
0.7156
$\pm$ 0.0002

### v2

@@ 1 Introduction
Despite the undeniable benefits of generative AI across various domains such as image synthesis and content creation [23, 40, 44, 20], growing concerns have emerged regarding their potential misuse and unintended consequences [45, 7, 46]. In particular, generative models can inadvertently memorize and reproduce identifiable faces or proprietary images from their training data, posing significant risks to intellectual property rights and personal privacy [32, 15, 5].
@@
These issues have urged the right to be forgotten [29, 41], a movement supporting individuals’ ability to request the removal of personal data (e.g., images of one’s face) from digital systems like AI models. While possible to remove requested images from training sets, retraining models from scratch is computationally impractical, especially if there are multiple requests over time [22, 38].
To address the growing need for post-hoc data removal, recent work has explored generative unlearning—the task of removing specific information from pretrained generative models without retraining from scratch [2, 42, 13, 30]. Approaches vary in how they suppress unwanted identities. Selective Amnesia [13] targets diffusion and VAE-based models, using continual learning techniques to push erased concepts into noise-like outputs. GUIDE [30], the first to address Generative Identity Unlearning (GUI) in GANs, instead manipulates the latent space by redirecting target embeddings toward a fixed mean identity representation. As illustrated in Figure 1, both methods demonstrate initial effectiveness for single-identity removal but face key limitations: (a) they lack support for forgetting multiple identities at once or over time; (b) their suppression mechanisms—mapping to noise or an average face—may leak information about the forgotten identity; and (c) they often degrade unrelated generations, causing semantic drift in retained identities. For example, in GUIDE, retained faces in the Baseline row of Figure 1 begin to converge toward the median face, reducing output diversity and model utility.
@@ 2 Related Works
Mitigating Misuse of Generative Models. As generative models continue to advance in their ability to synthesize and manipulate highly realistic images, concerns about ethical misuse have intensified. The integration of powerful image editing techniques—such as latent space manipulation [31], language-guided editing [26], and photorealistic face retouching [39]—has made it increasingly accessible for users to modify images of real individuals or apply stylistic transformations [33]. More alarmingly, recent studies have shown that these models can be exploited to extract memorized, copyrighted training data [3], often without the consent of the individuals or content owners involved. One of the most pressing risks is the unauthorized synthesis of a person’s likeness in misleading or harmful contexts, such as deepfakes or defamatory media [1, 24, 28]. To counter this, Thanh et al. [36] developed an adversarial attack against UNet, effectively corrupting generated images. Other works in the concept erasure domain focus on eliminating certain visual concepts. For example, Gandikota et al. [8] were among the first to introduce removing specific concepts from diffusion models using negative guidance. This is still an ongoing research direction with follow-up works such as [8, 25, 27] to better improve the balancing between unlearning efficiency and maintaining model utility on remaining concepts. Previous methods are identity-specific, requiring tailoring for each individual, which makes them non-scalable for protecting multiple identities and introduces substantial computational overhead. Moreover, these approaches primarily focus on associations between concepts, such as the “Van Gogh style,” and generated images, a focus more relevant to multi-modal learning, like text-to-image models. In contrast, our study centers on image-to-image models.
@@
Generative Unlearning. Machine unlearning is a technique that enables a machine learning model to “forget” or remove the influence of specific data points from its training data, without needing to be retrained from scratch [35, 2, 42]. The usage of approximate unlearning is more preferable in the context of GenAI because these models are often trained on large corpora and datasets, making the retraining approach infeasible [37, 21]. The interest in machine unlearning with generative models has grown recently, where the field was dominated by supervised unlearning before, i.e., classification tasks [12, 34, 10, 9, 4, 11]. Seo et al. [30] were the first to address GIU for GAN-based models in an approach called GUIDE. However, their method does not account for practical scenarios involving (i) simultaneous unlearning requests—an essential requirement in real-world applications where multiple individuals may request data removal at once—or (ii) sequential unlearning, where requests arrive at different times [43]. Our empirical results also demonstrate that while GUIDE is effective for unlearning, it hampers the model’s ability to maintain overall generation quality on retained identities.
@@
Recent studies have also addressed generative unlearning by mapping forgotten identities to a Gaussian distribution $\mathcal{N}(0,\sigma)$ while preserving the original distribution for remaining data in an approach called Selective Amnesia (SA) [13, 6]. However, using Gaussian noise to replace forgotten identities is suboptimal, as it fails to accurately represent certain training data distributions, potentially corrupting other identities. Furthermore, SA, as well as GUIDE, (iii) raise security concerns, as the association between a forgotten identity and its output (noise or average face) can still be detected. To this end, we introduce SUGAR to address each of these concerns (i, ii, and iii).
@@ 3.1 Problem Formulation
Problem formulation (multi-identity unlearning). We build on the GUI setup from Seo et al. [30] and extend it to handle multiple identity unlearning. We consider a pretrained EG3D-style pipeline $(E_{\psi},G_{\theta},R)$ and aim to obtain an unlearned generator $G_{\theta}^{u}$ that forgets a set of $K$ target identities while retaining performance on others. Let $\mathcal{D}=\mathcal{D}_{u}\cup\mathcal{D}_{r}$ denote the full dataset, where $\mathcal{D}_{u}$ contains images of the $K$ identities to remove and $\mathcal{D}_{r}$ the remainder ($|\mathcal{D}_{r}|\!\gg\!|\mathcal{D}_{u}|$). We operate in latent space: $W_{u}:=\{\,w=E_{\psi}(x)\mid x\in\mathcal{D}_{u}\,\}$ and $W_{r}:=\{\,w=E_{\psi}(x)\mid x\in\mathcal{D}_{r}\,\}$. We keep the mapping network and renderer $R$ fixed and update only $G$.
@@
Remapping function. To disassociate image generation from an identity $\mathcal{I}_{u}$ with latent vector $w_{u}$, a straightforward and effective strategy—used in GUIDE [30]—is to remap $w_{u}$ to a surrogate target identity $I^{avg}$ (e.g., the median face) with latent vector $w_{avg}$. With GUI, we adopt this idea and denote the GUIDE loss as the remapping objective.
@@ 3.2 ID-specific De-indentification
Unlike existing unlearning approaches with predefined target concepts, we propose a learnable and sample-specific procedure to determine where each forgotten sample should be mapped. We call this ID-specific De-identification.
Following Seo et al. [30], we define the identity latent vector $w_{id}$ for an individual as the difference between their latent representation $w_{u}$ and the average latent vector $\bar{w}$, i.e., $w_{id}=w_{u}-\bar{w}$. In our method, we train a mapping model $\mathcal{T}_{\Theta}$ to produce a new identity vector $w_{id}^{t}$ given any input latent vector $w_{u}\in\mathbb{R}^{d\times h\times w}$, such that:
@@
The intuition is that we learn a mapping $\mathcal{T}_{\Theta}$ that assigns each forgotten identity $w_{u}\in W_{u}$ a personalized counterfactual $w_{t}$: a different identity that (i) preserves coarse facial attributes of $w_{u}$ while (ii) being recognizably distinct by facial recognition metrics such as ID or LPIPS [30]. In latent space, $w_{t}$ serves as the surrogate for $w_{u}$; after unlearning, $G_{\theta}^{u}$ synthesizes faces consistent with $w_{t}$ rather than $w_{u}$. To learn this identity replacement, we train a transformation function $\Theta$ that maps each $w_{u}\in W_{u}$ to a replacement latent code $w_{t}=\Theta(\cdot,w_{u})$ that meets the above criteria. The transformation is supervised using the following objective:
@@ 3.3 Forgetting in a Continual Learning Fashion
Enforcing Identity Forgetting. Building on the remapping objective of Seo et al. [30] in Equation 1, we replace the fixed median-face target with a learnable de-identification target produced by our operator $\Theta(\cdot)$; that is, the average face $\mathcal{I}^{avg}$ is replaced by $\Theta(w)$. As noted in the GUIDE baseline, successful unlearning must also suppress identity information in the local neighborhood of each forgotten latent. We therefore sample a perturbed neighbor $w_{u,a}$ for each forgotten code $w_{u}$:
@@
Elastic Weight Consolidation (EWC). To efficiently preserve behavior in regions most susceptible to collateral damage, we adopt Elastic Weight Consolidation (EWC) [19]. We treat the vicinity probe set $\widetilde{W}$ as a proxy for prior behavior to be retained and regularize parameters that are important for agreement with the source model around forgotten latents.
@@ 4.1 Experimental Setup
Datasets. Following the settings of Seo et al. [30], we evaluated GUIDE in three forgetting scenarios: Random, InD (in-domain), and OOD (out-of-domain). These scenarios represent different sources of identities to unlearn. Specifically, identities are sampled randomly from the latent space (Random), from the FFHQ dataset [18] used for pre-training (InD), or from the CelebA-HQ dataset [17] as unlearning targets (OOD). For both in-distribution (InD) and out-of-distribution (OOD) settings, we use a GAN inversion network $\mathcal{I}$ to map each forgetting image $x_{u}^{i}\in\mathcal{D}_{u}$ to a latent code $w_{u}^{i}=E_{\psi}(x_{u}^{i})$, forming the forget set $W_{u}=\{w_{u}^{i}\}_{i=1}^{N}$; the retaining set $W_{r}$ is sampled from the same datasets and inverted analogously.
Baselines. We select GUIDE [30] as our baseline for comparison, as it aligns with our focus on identity-specific unlearning. GUIDE effectively removes identity-related features while preserving the overall structure and quality of the generated images. In contrast, other methods, such as Selective Amnesia (SA) [13] and Feng et al. [6], map forgotten identities to Gaussian noise. This approach often results in unnatural artifacts and disrupts facial reconstruction, making it unsuitable for our task. Specifically, mapping to Gaussian noise prevents the generation of coherent and realistic reconstructions after unlearning, which is a critical requirement in our setting.
Evaluation Metrics. We use Identity Similarity (ID), computed with the face recognition network CurricularFace [16], to measure the similarity between images of faces generated from the same latent codes before and after unlearning. Additionally, we use the Fréchet Inception Distance (FID) score [14] to quantify the overall visual quality and distribution alignment of the generated images. These metrics are calculated on both the forgetting set $\mathcal{D}_{u}$ and the retaining set $\mathcal{D}_{r}$, where $\mathcal{D}_{r}$ is sampled from the same datasets to which $\mathcal{D}_{u}$ belongs.
@@ 4.2.1 Balancing Forgetting and Retaining
#IDS Methods FFHQ (In-domain Distribution) CelebAHQ (OOD Distribution) Random
@@
ID GUIDE 0.2773       $\pm$0.0099 0.0040       $\pm$0.0077 0.0095       $\pm$0.0017 -0.0275       $\pm$0.0034 -0.0431       $\pm$0.0018
Ours 0.3576       $\pm$0.0146 0.2664       $\pm$0.0134 0.2559       $\pm$0.0078 0.2917       $\pm$0.0050 0.2433       $\pm$0.0051
FID GUIDE 133.90       $\pm$6.9372 177.27       $\pm$2.2242 163.83       $\pm$0.9667 143.44       $\pm$0.8456 146.43       $\pm$0.6706
Ours 115.98       $\pm$7.7846 125.98       $\pm$0.8796 125.93       $\pm$0.3963 105.40       $\pm$0.6074 106.12       $\pm$1.4283
@@ 5 Ablation Study
$d$-Controllable Unlearning. Our method can be used as a controllable unlearning solution in I2I generative models, where $d$ is the control coefficient. In Equation 3, with a well-trained model $\Phi$, the targeted identity for each identity $w_{id}$ is adjusted with a different coefficient $d$. As shown in Figure 6, the higher the value of $d$, the more features of the forgetting ID $w_{id}$ are retained in the new mapping identity $w_{t}$. With our method, the model vendor can select this controllable parameter to balance the trade-off and the forgetting level that reaches a consensus between the vendor and the requested unlearning party. However, there is a trade-off between forgetting and retaining utility, as discussed in previous work [6, 42]. This relationship is presented in Table 3. While reducing the distance can make the model forget the forgetting set $\mathcal{D}_{f}$ more drastically (i.e., lower ID and higher FID for the forgetting set), this leads to a greater impact on the retaining set (i.e., lower ID). It is worth noting that even with $d=30$, the generated images are much different from the original identities.
@@
Dual Objective: Identity Removal and Global Utility Preservation. We study the impact of our two loss terms: forgetting neighbor loss $\mathcal{L}_{nei}$ for thorough identity forgetting and the global preservation loss $\mathcal{L}_{\text{ewc}}$ for utility retention. To evaluate identity removal, we conduct a multi-image test on CelebAHQ identities, assessing whether our method forgets other images of the same identity that are not explicitly included in $W_{u}$. Figure 9 summarizes the results. “Ours-v1” denotes the variant without the neighbor loss $\mathcal{L}_{nei}$ in Equation 5, so the forgetting objective applies only to the anchor images of each target identity. Without this term, residual attributes of the forgotten identity can persist, even when the generated images depict distinct individuals. Adding $\mathcal{L}_{nei}$ enforces consistency across the local neighborhood in latent space, which is crucial for the ultimate goal of removing an entire identity rather than only a single image. To analyze $\mathcal{L}_{ewc}$, Figure 9 compares reconstructions with (Ours) and without (Ours-v2) the global preservation loss. The first two retained identities come from the same distribution as the forgotten ones, while the last three (IDs 3–5) are randomly sampled latents. Random identities remain largely unaffected by unlearning (even under GUIDE), whereas closer identities (IDs 1–2) are more susceptible to collateral forgetting. Ours-v2 already preserves these better than GUIDE due to our de-identification, and adding $\mathcal{L}_{ewc}$ further improves retention by preserving fine details (e.g., glasses, hair, wrinkles). We provide a detailed quantitative result for our ablation study in Appendices B.1. and B.2.
Sequential unlearning. We consider the setting where unlearning requests arrive sequentially rather than in a single batch. Let $G_{\theta}^{(0)}\!=\!G_{\theta}^{*}$ be the pretrained model and let $\mathcal{D}_{f}^{(k)}$ denote the (new) identities to forget at stage $k{=}1,\dots,4$. At each stage, we initialize from the previously unlearned model and run the same unlearning procedure (architecture, losses, schedule, and hyperparameters unchanged): $G_{\theta}^{(k)}\;\leftarrow\;\textsc{Unlearn}\!\big(G_{\theta}^{(k-1)},\,\mathcal{D}_{f}^{(k)}\big).$ Figure 9 shows qualitative results after four stages for GUIDE and SUGAR. While GUIDE accumulates collateral damage—progressively eroding features of the retained identities across stages—SUGAR continues to forget the requested identities while preserving the appearance of the retained set. Detailed setup and additional analysis are provided in Appendix B.5.
@@
Forgetting GUIDE 254.31 $\pm$ 376.36 25.4262 $\pm$ 0.2026 0.2048 $\pm$ 0.0000 25.4262 $\pm$ 0.2026
Ours 1022.81 55301.1 18.5459 $\pm$ 0.8418 0.2936 $\pm$ 0.0002 0.7156 $\pm$ 0.0002